OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (-2, -3) and has a slope of 1/2? Answer 2x + y = 4 2x - y = 4 x + 2y = 4 x - 2y = 4

OpenStudy (anonymous):

Answer is either B or D

Parth (parthkohli):

I'd use point-slope form first...\[\rm y - y_1 = m(x - x_1)\]then convert into standard form.

Parth (parthkohli):

Substituting the information,\[\rm y - (-3) = 0.5(x - (-2))\]or\[\rm y + 3 = 0.5(x + 2)\]

OpenStudy (anonymous):

:) Correct

Parth (parthkohli):

Multiply both sides by 2.\[\rm 2y + 6 = x + 2\]

Parth (parthkohli):

Yup :D

Parth (parthkohli):

Now bring the variables on the left side and the constants on the right.

OpenStudy (anonymous):

So far so good